2 A4 in schedule 40steel pipe(k20 wm-K) carrying steam at 200t C, is placed in air at 30 The convective film coefficients between the air and the pipe outer surface_and between steam and pipe inner surface are 4 and 500 wm2-C respectively. It is desired to reduce the heat loss to 60 % of its present-value by the addition ofan insulation with k-04 wm- C. Assuming that the pipe surface temperature and the exposed surface film coefficient remain unchanged as the insulation is added, find the required thickness of insulation in.cm,(15 p
Answer #1

Given :

4 in Schedule 40 steel pipe

from internet sources you can find details like wall thickness of pipe is 6.02 mm.

k = 20 w/m-K

steam temp. 200oC

air temp = 30oC

Convective film coefficient between air & the pipe outer surface = 4 w/m2-oC

Convective film coefficient between steam & the pipe inner surface = 500 w/m2-oC

we want to reduce the heat loss to 60 % of its present value .

By adding insulation of material with k = 0.4 w/m-oC

to find required thickness of insulation in cm.

heat loss when insulation was not provided

Q=UA(T2-T1)

1/U=1/hi+l/k+1/ho

1/U =1/4+0.00602/20+1/500

1/U= 0.252301

U=3.9635 w/m2-oC

Q/A=3.9635(200-30) = 673.795 w/m2

we need to reduce loss to its 60%

desired loss = 0.6 *673.795 =404.277 w/m2

now

404.277 =U*\DeltaT= U*(200-30)

\thereforeU =2.3781 w/m2-oC

where after adding insulation

let required thickness is t

1/U =1/hi+l/k+1/ho+t/ki

0.4205 = 1/4+0.00602/20+1/500+t/0.4

0.4205 = 0.252301+2.5t

2.5t =0.168199

t =0.0672796 m

t=6.72 cm